OpenStudy (anonymous):

I WILL MEDAL AND FAN!! PLEASE! Which of the following best describes the function f(x) + g(x)? x 1 2 3 4 5 f(x) 4 16 64 256 1,024 g(x) -1 -1 -1 -1 -1 A. 4^x - 1x B.-1(4^x) C. 4^x^-1 D. 4^x - 1

OpenStudy (anonymous):

Omg please someone!

OpenStudy (anonymous):

I really need to submit this test!!

OpenStudy (anonymous):

Let's start with the easy one. What is g(x)?

OpenStudy (anonymous):

When you plug in the x value for for the equations in your answer choice you should see same answers for example if you plug in x=1 you should get 4^1-1(1)=4+(-1)

OpenStudy (anonymous):

you would have to plug in all values of x to get a definitive answer

OpenStudy (anonymous):

4^x-1x would be the equation from the answer choice and 4+(-1)would be f(x)+g(x)

OpenStudy (anonymous):

Oh ok I think I get it.. What is a definitive answer tho? @arturos

OpenStudy (anonymous):

like you are 100% sure that the answer choice you choose is the correct one

OpenStudy (anonymous):

Oh ok... So when we plug in all values of x... We get the 3rd or the 4th answer?

OpenStudy (anonymous):

@arturos

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But im so confused whats the difference between 4^x^-1 which is the 3rd answer and 4^x -1

OpenStudy (anonymous):

@arturos

OpenStudy (anonymous):

@Vandreigan I dont understand what you mean by that?

OpenStudy (anonymous):

oh -1?

OpenStudy (anonymous):

Yeah. g(x) = -1. It's a constant. Can you see what f(x) is?

OpenStudy (anonymous):

\[4^{x ^{-1}}\] is 4^x^-1 where 4^x-1 is\[4^{x}-1\]

OpenStudy (anonymous):

oh ok thank you! @arturos let me just see what @Vandreigan was going to say

OpenStudy (anonymous):

Yeah no problem

OpenStudy (anonymous):

Well it is exponential right? it 4 times 2 and then 4 times3 and so o right?

OpenStudy (anonymous):

Not quite. The series is: 4, 16, 64, 256, 1024 16 isn't 4*2, but you're not too far off.

OpenStudy (anonymous):

Re-reading, I think you meant 4^2, not 4*2, which is correct. so f(x) is what?

OpenStudy (anonymous):

16 is 4^2

OpenStudy (anonymous):

Oh yes that is what i meant srry...

OpenStudy (anonymous):

Yep, so when x = 1, we get 4^1. When x=2, we have 4^2. Etc. So what is f(x)?

OpenStudy (anonymous):

4^x?

OpenStudy (anonymous):

Yep :) so, f(x) = 4^x g(x) = -1 So what is f(x) + g(x)?

OpenStudy (anonymous):

Ohhh! Wow thanks that makes so much sense when you put it like that! So its would 4^x -1

OpenStudy (anonymous):

You got it :)

OpenStudy (anonymous):

or the last one right now?

OpenStudy (anonymous):

Yep. D.

OpenStudy (anonymous):

Well done :)

OpenStudy (anonymous):

im sorry if i wasn't explaining clearly enough just let me know and ill try to work on my explanations

OpenStudy (anonymous):

No its ok im just better with wording rather than just equations